# SoK: A Classification for AI-driven Personalized Privacy Assistants

### v2

@@ B.1. Threat I – Planning Limitations of the SLR
The first threat relates to the planning of the SLR in terms of identifying the need and justification for this study. Here, we were concerned with identifying existing reviews (systematic and non-systematic) on the topic of AI-driven PPAs. The initial searches did not reveal any review studies on the topic, as described in Section 3.1, pointing to a significant gap in secondary research AI PPAs. The planning phase of the SLR is also critical to outline the research questions and provide the basis for an objective investigation of the studies that are being reviewed. If the RQs are not explicitly stated or omit the key topics, the literature review results can be flawed, overlooking the key information. To mitigate this threat, we outlined two RQs and objectives (Section 3.1). In summary, we seek to minimize any bias or limitations during the planning phase when defining the scope and objectives of this SLR. As a last step in the planning phase, the team finalized and cross-checked the study protocol to minimize the limitations of the SLR plan before proceeding to the subsequent phases.
B.2. Threat II – Validity of the Search Process
@@
Identifying and selecting the studies reviewed in the SLR are also significant processes to be observed. Selecting studies is a critical step; if any relevant papers are missed, the results of the SLR may be flawed. Therefore, we followed a stepwise process (Section 3.2), starting with a literature screening and followed by a complete reading of papers. This selection process was carried out independently by two reviewers. We also performed forward and backward snowballing, looking for references to other potentially relevant studies. Also, this SLR restricts the selection of publications to four scientific databases: Scopus, Web of Science, IEEE Xplore, and ACM Digital Library. These databases were used due to their high relevancy to computer science, privacy, and data protection, as well as to maintain a feasible search space. This search process gives us confidence that we minimized limitations related to (i) excluding or overlooking relevant studies or (ii) including irrelevant studies that could impact the results and their reporting in the SLR.
B.3. Threat III – Potential Bias in the Synthesis Process
@@
Some threats should also be considered regarding the potential bias in synthesizing the data from the review and documenting the results. This means that if there are some limitations in the data synthesis, they directly impact the results of this SLR. Typical examples of such limitations could be a flawed research taxonomy and a mismatch of potential research gaps. To minimize the bias in synthesizing and reporting the results, we have created a data extraction form that uses well-known classification schemes, such as the ones proposed by Wieringa et al. and Creswell and Creswell [2017], or Arrieta et al. for the classification of AI. Three researchers independently reviewed this data extraction form while revising the research protocol. While one of the researchers led the data extraction step, two other authors helped by cross-checking the work throughout the process for consistency. Three authors were involved in the creation of the classification scheme derived from the literature (i.e., shown in Section 2), actively working on reviewing the list of categories for consistency through a series of meetings. Furthermore, this SLR also offers a complete replication package, enabling researchers to reproduce or extend this review (https://anonymous.4open.science/r/SoK_AI_PPA-E29F).

### v4

@@ VII-A Threat I – Planning Limitations of the SLR
The first threat relates to the planning of the SLR in terms of identifying the need and justification for this study. Here, we were concerned with identifying existing reviews (systematic and non-systematic) on the topic of AI-driven PPAs. The initial searches did not reveal any review studies on the topic, as described in Section III-A, pointing to a significant gap in secondary research AI PPAs. The planning phase of the SLR is also critical to outline the research questions and provide the basis for an objective investigation of the studies that are being reviewed. If the RQs are not explicitly stated or omit the key topics, the literature review results can be flawed, overlooking the key information. To mitigate this threat, we outlined two RQs and objectives (Section III-A). In summary, we seek to minimize any bias or limitations during the planning phase when defining the scope and objectives of this SLR. As a last step in the planning phase, the team finalized and cross-checked the study protocol to minimize the limitations of the SLR plan before proceeding to the subsequent phases.
VII-B Threat II – Validity of the Search Process
@@
Identifying and selecting the studies reviewed in the SLR are also significant processes to be observed. Selecting studies is a critical step; if any relevant papers are missed, the results of the SLR may be flawed. Therefore, we followed a stepwise process (Section III-B), starting with a literature screening and followed by a complete reading of papers. This selection process was carried out independently by two reviewers. We also performed forward and backward snowballing, looking for references to other potentially relevant studies. Also, this SLR restricts the selection of publications to four scientific databases: Scopus, Web of Science, IEEE Xplore, and ACM Digital Library. These databases were used due to their high relevancy to computer science, privacy, and data protection, as well as to maintain a feasible search space. This search process gives us confidence that we minimized limitations related to (i) excluding or overlooking relevant studies or (ii) including irrelevant studies that could impact the results and their reporting in the SLR.
VII-C Threat III – Potential Bias in the Synthesis Process
@@
Some threats should also be considered regarding the potential bias in synthesizing the data from the review and documenting the results. This means that if there are some limitations in the data synthesis, they directly impact the results of this SLR. Typical examples of such limitations could be a flawed research taxonomy and a mismatch of potential research gaps. To minimize the bias in synthesizing and reporting the results, we have created a data extraction form that uses well-known classification schemes, such as the ones proposed by Wieringa et al. [64] and Creswell and Creswell [81], or Arrieta et al. [13] for the classification of AI. Three researchers independently reviewed this data extraction form while revising the research protocol. While one of the researchers led the data extraction step, two other authors helped by cross-checking the work throughout the process for consistency. Three authors were involved in the creation of the classification scheme derived from the literature (i.e., shown in Section 2), actively working on reviewing the list of categories for consistency through a series of meetings. Furthermore, this SLR also offers a complete replication package, enabling researchers to reproduce or extend this review (https://github.com/Victor-Morel/SLR_AI_PPA).
VIII Conclusion
With the SLR presented in this article, we provide a classification and common vocabulary to compare and discuss AI-driven PPAs. Although many papers (41 in our selection) have already been published on AI-driven PPAs in the last decade, they do not yet form a coherent body of literature. AI-driven PPAs can be improved by performing standard evaluations (including their usability), integrating privacy by design in their design process, providing additional explanations for their decisions, and considering a broader range of system context and larger variety of data sources. We hope this survey and its classification allow users and developers of AI-driven PPAs to compare different solutions and understand their pros and cons. Moreover, the recommendations should help improve AI-driven PPAs in different ways, addressing the challenges raised by AI’s latest developments (including LLMs), data collection, and modern regulations.